A stone is thrown horizontally with the velocity of 15 m/s. Determine the tangential and normal accelerations of the stone in one second after it begins to move

Grade:11

2 Answers

Vikas TU
7 years ago
In x direcn.,
Vx = 15cos(thetha)
In y direcn.,
Vy = 15sin(thetha) – g*1
 
acclern. in x dircn. = > (Vx – 15costhetha)/1 = 0
accln. in y dircn. = > Vy – 15sinthetha)/1 = > -g
Hence net acclen.is in downward direcn that is => g m/s^2
Sonia
6 years ago
At=2g/√13 an=3g/√13 ßhhsejuxjshuxxeiis Ed uecudjd it up on my way back to you as well as well and that the message and all of fruit